# Gorenstein singularities with $\mathbb{G}_{m}$-action and moduli spaces of holomorphic differentials

### v1

@@ 5.9. The stratum $\mathbb{P}\mathcal{H}(6,2)^{\operatorname{odd}}$
[…]
@@
Since $x$, $y$, $z$, $w$, and $u$ have weights $14$, $21$, $12$, $15$, and $18$ under the $\mathbb{G}_{m}$-action, respectively, the standard projectivization of the corresponding Gorenstein curve $Y$ as well as its deformations $(C,p_{1},p_{2})\in\mathbb{P}\mathcal{H}(6,2)^{\operatorname{odd}}$ can be embedded in the weighted projective space $\mathbb{P}(14,21,12,15,18,1)$, where the divisor at infinity is cut out by the hyperplane of weight $1$ as $\frac{1}{3}p_{1}+\frac{1}{7}p_{2}$.
@@
The nontrivial weights are $3$, $6$, $7$, $9$, and $21$. We thus obtain that
@@ 5.10. The stratum $\mathbb{P}\mathcal{H}(5,3)$
[…]
@@
Since $x$, $y$, $z$, and $w$ have weights $6$, $9$, $8$, and $10$ under the $\mathbb{G}_{m}$-action, respectively, the standard projectivization of the corresponding Gorenstein curve $Y$ as well as its deformations $(C,p_{1},p_{2})\in\mathbb{P}\mathcal{H}(5,3)$ can be embedded in the weighted projective space $\mathbb{P}(6,9,8,10,1)$, where the divisor at infinity is cut out by the hyperplane of weight $1$ as $\frac{1}{2}p_{1}+\frac{1}{3}p_{2}$.
@@
The nontrivial weights are $2$, $3$, $4$, $6$, and $12$. We thus obtain that
@@ 5.11. The varying stratum $\mathbb{P}\mathcal{H}(2,2,2)^{\operatorname{even}}$
As mentioned in Remark 3.1, it was already observed in [Bat24, Remark 2.1] that the singularity classes corresponding to the varying stratum $\mathbb{P}\mathcal{H}(1,1,1,1)$ in genus three have one-dimensional moduli, determined by the cross-ratio of the four zeros in the line section of the canonical map. For completeness, here we consider another varying stratum $\mathbb{P}\mathcal{H}(2,2,2)^{\operatorname{even}}$ in genus four. We will show that its singularity classes are not unique.
In this case, we have $g=4$, $b=3$, and $\delta=6$. Recall the discussion in Section 5.8, which implies that the gap sequence for this even spin component is [2 1 $\ 1$][ 2 1 1 ].
@@
    First, suppose $t_{1}\oplus t_{2}\oplus t_{3}\in R$. Since each individual $t_{i}^{3}\not\in R$, up to reordering and rescaling the parameters, we can assume that a quadratic generator is $t_{1}^{2}\oplus at_{2}^{2}\oplus 0$, where $a\neq 0,1$. Up to rescaling the parameters, we can assume that $a=-1$.
@@
    $y$ $=t_{1}^{2}\oplus(-t_{2}^{2})\oplus 0.$
    $y(y-x^{2})(y+x^{2}).$
@@
    $(xy,y(z^{2}-y^{3}),xz(z-x^{3}),z^{2}-y^{3}-x^{3}z).$
    Since $x$, $y$, and $z$ have weights $1$, $2$, and $3$ under the $\mathbb{G}_{m}$-action, respectively, the standard projectivization of the corresponding Gorenstein curve $Y$ as well as its deformations $(C,p_{1},p_{2},p_{3})$ can be embedded in the weighted projective space $\mathbb{P}(1,2,3,1)$, where the divisor at infinity is cut out by the hyperplane of weight $1$ as $p_{1}+p_{2}+p_{3}$.
@@
The nontrivial weights are $1$, $1$, $2$, and $3$. We thus obtain that
@@ 5.12. Nonvarying strata with ordinary marked points
We can generalize the definition of nonvarying strata by including zeros of order 00, i.e., ordinary marked points. Let $\mu^{\prime}=(m_{1},\ldots,m_{n},0,\ldots,0)$, where $\mu=(m_{1},\ldots,m_{n})$ is one of the above nonvarying signatures and the new parameters $s_{1},\ldots,s_{k}$ correspond to the branches of the last $k$ ordinary points. Up to rescaling the parameters, we can assume that
@@
$y$ $=t_{1}^{3}\oplus 0\oplus 0\cdots\oplus 0,$
@@
5.13. Nonvarying strata without obstructions
@@
$\mu=$ $(3,1),(2,2)^{\operatorname{odd}},(2,1,1),(5,1),(4,2)^{{\rm even}/{\rm odd}},(3,3)^{\operatorname{nonhyp}},(3,2,1),(5,3)$
@@
As mentioned before, we can verify that $\dim T^{2}=0$ for these strata by using Singular. Alternatively, for each of the strata listed above, the defining ideal of the corresponding singularity requires at most four parameters, and Gorenstein singularities with embedding codimension at most three have no infinitesimal obstructions; see [Her80]. Consequently, $\mathbb{P}\mathcal{H}(\mu)=\mathbb{P}{\rm Def}(Y,q)^{-}_{s}$ is an open subset of the weighted projective space $\mathbb{P}(T^{1,-})$; see Section 2.1. Note that these nonvarying strata are known to be affine; see [Che24, Theorem 1.1]. It implies that the complement of each $\mathbb{P}\mathcal{H}(\mu)$ in its $\mathbb{P}(T^{1,-})$ is a hypersurface of pure codimension $1$; see [Goo69, Proposition 1], where the hypersurface is the discriminant locus parameterizing singular deformations in $\mathbb{P}{\rm Def}(Y,q)^{-}$. Regarding the explicit description of the corresponding weighted projective space, see Example 5.2 below for a demonstration. The remaining claims thus follow from the facts that weighted projective spaces are rational varieties and the rational Chow ring of a weighted projective space is generated by the class of a hypersurface; see [Inc22, Lemma 4.7]. ∎

### v2

@@ 5.9. The stratum $\mathbb{P}\mathcal{H}(6,2)^{\operatorname{odd}}$
In this case, we have $g=5$, $b=2$, and $\delta=6$. The gap sequence is determined by $\alpha_{1}\ \alpha_{2}\ \alpha_{3}\ \alpha_{4}\ \alpha_{5}\ \alpha_{6}\ 1]$, where $\alpha_{7}=1$ is the last nonzero entry and $\sum_{i\geq 1}\alpha_{i}=5$. Since here we consider the odd spin component, $\dim H^{0}(Y,3y_{1}+y_{2})=1$, and hence any $t_{1}^{n_{1}}\oplus t_{2}^{n_{2}}\in R$ with $n_{1}\leq 3$ and $n_{2}\leq 1$ must be the constant function. Note that the exponents of nonzero terms of $t_{1}$ and $t_{2}$ in every (non-decomposable) generator of $\mathfrak{m}/\mathfrak{m}^{2}$ are proportional as $7:3$. Hence, $t_{1}^{i}\oplus 0\not\in R$ for $i=1,2,3$ and $0\oplus t_{2}\not\in R$. It implies that the gap sequence is [2 1 1 0 0 0 1][2\ 1\ 1\ 0\ 0\ 0\ 1].
@@
Since $x$, $y$, $z$, $w$, and $u$ have weights $14$, 2121, $12$, $15$, and $18$ under the $\mathbb{G}_{m}$-action, respectively, the standard projectivization of the corresponding Gorenstein curve $Y$ as well as its deformations $(C,p_{1},p_{2})\in\mathbb{P}\mathcal{H}(6,2)^{\operatorname{odd}}$ can be embedded in the weighted projective space $\mathbb{P}(14,21,12,15,18,1)$, where the divisor at infinity is cut out by the hyperplane of weight $1$ as $\frac{1}{3}p_{1}+\frac{1}{7}p_{2}$.
@@
The nontrivial weights are $3$, 66, 77, 99, and $21$. We thus obtain that
@@ 5.10. The stratum $\mathbb{P}\mathcal{H}(5,3)$
In this case, we have $g=5$, $b=2$, and $\delta=6$. The gap sequence is determined by $\alpha_{1}\ \alpha_{2}\ \alpha_{3}\ \alpha_{4}\ \alpha_{5}\ 1]$, where $\alpha_{6}=1$ is the last nonzero entry and $\sum_{i\geq 1}\alpha_{i}=5$. Since $\alpha_{6}\neq 0$, we have $\alpha_{1},\alpha_{2},\alpha_{3}\neq 0$. Moreover, the exponents of nonzero terms of $t_{1}$ and $t_{2}$ in every (non-decomposable) generator of $\mathfrak{m}/\mathfrak{m}^{2}$ are proportional as $6:4=3:2$. Therefore, $\alpha_{1}=2$, and hence the gap sequence is [2 1 1 0 0 1][2\ 1\ 1\ 0\ 0\ 1].
@@
Since $x$, $y$, $z$, and $w$ have weights $6$, 99, 88, and $10$ under the $\mathbb{G}_{m}$-action, respectively, the standard projectivization of the corresponding Gorenstein curve $Y$ as well as its deformations $(C,p_{1},p_{2})\in\mathbb{P}\mathcal{H}(5,3)$ can be embedded in the weighted projective space $\mathbb{P}(6,9,8,10,1)$, where the divisor at infinity is cut out by the hyperplane of weight $1$ as $\frac{1}{2}p_{1}+\frac{1}{3}p_{2}$.
@@
The nontrivial weights are $2$, 33, 44, 66, and $12$. We thus obtain that
@@
5.11. The varying stratum $\mathbb{P}\mathcal{H}(1,1,1,1)$
As mentioned in Remark 3.1, it was already observed in [Bat24, Remark 2.1] that the singularity classes corresponding to the varying stratum $\mathbb{P}\mathcal{H}(1,1,1,1)$ in genus three have one-dimensional moduli, determined by the cross-ratio of the four zeros in the line section of the canonical map. There is a special case when the underlying curve is hyperelliptic and the four zeros form two pairs of hyperelliptic conjugate points. In what follows we analyze the two cases separately.
  1. (1)
    First, the canonical model of a non-hyperelliptic smooth curve $C$ of genus three is a plane quartic, where a canonical divisor $p_{1}+p_{2}+p_{3}+p_{4}$ in $C$ is cut out by a line $L$. The resulting singularity $(Y,q)$ corresponds to a cone of four lines through $p_{1},p_{2},p_{3}$, and $p_{4}$ that meet at $q$. Without loss of generality, we can write down the defining equation of $(Y,q)$ as xy (x-y)(x- cy), where $c\neq 0,1,\infty$ is determined by the cross-ration of the four zeros in $L$.
    Since in this case $(Y,q)$ is a planar curve singularity, its miniversal deformation space can be described by
    $xy(x-y)(x-cy)+\lambda_{1}x^{2}y^{2}+\lambda_{2}x^{2}y+\lambda_{3}xy^{2}$
[…]
    Since $x$ and $y$ both have weight $1$, the weights of the deformation parameters $\lambda_{1},\ldots,\lambda_{9}$ are $0,1,1,2,2,2,3,3,4$, respectively. In particular, $\mathbb{P}{\rm Def}^{-}(Y,q)$ is isomorphic to the weighted projective space $\mathbb{P}(1,1,2,2,2,3,3,4)$. Additionally, the deformation space $\mathbb{P}{\rm Def}^{0}(Y,q)$ of weight 0 given by the direction of $\lambda_{1}$ parameterizes equisingular deformations that alter the cross-ratio of the four zeros in $L$.
  2. (2)
    Next, suppose $(Y,q)$ arises from the degeneration of a hyperelliptic curve $C$. Up to reordering the zeros, we can assume that $p_{1}+p_{2}\sim p_{3}+p_{4}$ is the hyperelliptic $g^{1}_{2}$. Let $t_{i}$ be the parameter of the rational branch attached to each $p_{i}$. Up to rescaling the parameters,
    $x$ $=t_{1}\oplus{\rm i}t_{2}\oplus 0\oplus 0,$
    $y$ $=0\oplus 0\oplus t_{3}\oplus{\rm i}t_{4},$
    $z$ $=t_{1}^{2}\oplus 0\oplus(-t_{3}^{2})\oplus 0$
    generate the local ring of the singularity $(Y,q)$. In this case, the defining equations of $(Y,q)$ are (xy, z(z - x^2 - y^2)). Note that it is not a planar singularity.
    By using the software Singular, we can describe the miniversal deformation space of this hyperelliptic singularity as follows:
    $\begin{pmatrix}xy\\ z^{2}-x^{2}z-y^{2}z\\ \end{pmatrix}+\lambda_{1}\begin{pmatrix}z\\ 0\\ \end{pmatrix}+\lambda_{2}\begin{pmatrix}y\\ 0\\ \end{pmatrix}+\lambda_{3}\begin{pmatrix}x\\ 0\\ \end{pmatrix}+\lambda_{4}\begin{pmatrix}1\\ 0\\ \end{pmatrix}$
[…]
    Since the weights of $x$, $y$, and $z$ are $1$, 11, and $2$, respectively, the weights of the deformation parameters $\lambda_{1},\ldots,\lambda_{9}$ are $0,1,1,2,2,2,3,3,4$, respectively. Interestingly, $\mathbb{P}{\rm Def}^{-}(Y,q)$ is still isomorphic to the weighted projective space $\mathbb{P}(1,1,2,2,2,3,3,4)$. Additionally, the deformation space $\mathbb{P}{\rm Def}^{0}(Y,q)$ of weight 0 given by the direction of $\lambda_{1}$ deforms this hyperelliptic singularity back to the preceding case of planar singularities by using the relation $xy+\lambda_{1}z=0$.
In summary, the union of $\mathbb{P}{\rm Def}^{-}(Y,q)$ for $(Y,q)\in S(1,1,1,1)$ forms a weighted projective bundle with fiber $\mathbb{P}(1,1,2,2,2,3,3,4)$ over the one-dimensional base of cross-ratios, where the degenerate cross-ratios correspond to the three hyperelliptic loci given by $p_{1}+p_{2}\sim p_{3}+p_{4}$, $p_{1}+p_{3}\sim p_{2}+p_{4}$, or $p_{1}+p_{4}\sim p_{2}+p_{3}$. In particular, $\mathbb{P}\mathcal{H}(1,1,1,1)$ is the complement of the discriminant locus of singular deformations in this weighted projective bundle.
@@ 5.12. The varying stratum $\mathbb{P}\mathcal{H}(2,2,2)^{\operatorname{even}}$
For completeness, here we consider another varying stratum $\mathbb{P}\mathcal{H}(2,2,2)^{\operatorname{even}}$ in genus four. We will show that its singularity classes are not unique.
In this case, we have $g=4$, $b=3$, and $\delta=6$. Recall the discussion in Section 5.8, which implies that the gap sequence for this even spin component is [2 1 1][2\ 1\ 1].
@@
    First, suppose $t_{1}\oplus t_{2}\oplus t_{3}\in R$. Since each individual $t_{i}^{3}\not\in R$, up to reordering and rescaling the parameters, we can assume that a quadratic generator is $t_{1}^{2}\oplus at_{2}^{2}\oplus 0$, where $a\neq 0,1$.
@@
    $y$ $=t_{1}^{2}\oplus(at_{2}^{2})\oplus 0.$
    In terms of these generators, the defining equation of this singularity is y(y-x^2)(y- ax^2).
@@
    In terms of these generators, the defining equations of this singularity are (xy, z^2 - y^3 - x^3z).
    Since $x$, $y$, and $z$ have weights $1$, 22, and $3$ under the $\mathbb{G}_{m}$-action, respectively, the standard projectivization of the corresponding Gorenstein curve $Y$ as well as its deformations $(C,p_{1},p_{2},p_{3})$ can be embedded in the weighted projective space $\mathbb{P}(1,2,3,1)$, where the divisor at infinity is cut out by the hyperplane of weight $1$ as $p_{1}+p_{2}+p_{3}$.
@@
The nontrivial weights are $1$, 11, 22, and $3$. We thus obtain that
@@
5.13. Nonvarying strata without obstructions
Theorem 5.1.
$\mu=\$$(3,1),(2,2)^{\operatorname{odd}},(2,1,1),(5,1),(4,2)^{{\rm even}/{\rm odd}},(3,3)^{\operatorname{nonhyp}},(3,2,1),(5,3)$
Proof.
As mentioned before, we can verify that $\dim T^{2}=0$ for these strata by using Singular. Alternatively, for each of the strata listed above, the defining ideal of the corresponding singularity requires at most four parameters, and Gorenstein singularities with embedding codimension at most three have no infinitesimal obstructions; see [Her80]. Consequently, $\mathbb{P}\mathcal{H}(\mu)=\mathbb{P}{\rm Def}^{-}_{s}(Y,q)$ is an open subset of the weighted projective space $\mathbb{P}(T^{1,-})$; see Section 2.1. Note that these nonvarying strata are known to be affine; see [Che24, Theorem 1.1]. It implies that the complement of each $\mathbb{P}\mathcal{H}(\mu)$ in its $\mathbb{P}(T^{1,-})$ is a hypersurface of pure codimension $1$; see [Goo69, Proposition 1], where the hypersurface is the discriminant locus parameterizing singular deformations in $\mathbb{P}{\rm Def}^{-}(Y,q)$. Regarding the explicit description of the corresponding weighted projective space, see Example 5.2 below for a demonstration. The remaining claims thus follow from the facts that weighted projective spaces are rational varieties and the rational Chow ring of a weighted projective space is generated by the class of a hypersurface; see [Inc22, Lemma 4.7]. ∎
Example 5.2.
Consider the nonvarying stratum $\mathbb{P}\mathcal{H}(2,1,1)$. The defining equations of the corresponding singularity $(Y,q)$ have been given in Section 5.3. With the defining equations at hand, we can compute its miniversal deformation space by using the software Singular as follows:
  • LIB "sing.lib";
  • ring R = 0,(x,y,z),dp;
  • ideal I = xy, z2-y3-x2z;
  • deform(I);
$\begin{pmatrix}xy\\ z^{2}-y^{3}-x^{2}z\\ \end{pmatrix}$ $+\lambda_{1}\begin{pmatrix}0\\ z\\ \end{pmatrix}+\lambda_{2}\begin{pmatrix}0\\ y\\ \end{pmatrix}+\lambda_{3}\begin{pmatrix}0\\ x\\ \end{pmatrix}+\lambda_{4}\begin{pmatrix}0\\ 1\\ \end{pmatrix}$
$+\lambda_{5}\begin{pmatrix}z\\ 0\\ \end{pmatrix}+\lambda_{6}\begin{pmatrix}y\\ 0\\ \end{pmatrix}+\lambda_{7}\begin{pmatrix}x\\ 0\\ \end{pmatrix}+\lambda_{8}\begin{pmatrix}1\\ 0\\ \end{pmatrix}$.
The weights of $x$, $y$, and $z$ are $3$, 44, and $6$, respectively. The weights of the deformation parameters $(\lambda_{1},\ldots,\lambda_{8})$ can be read off from those of $(z,y^{2},xz,y^{3},xy/z,x,y,xy)$ which are $(6,8,9,12,1,3,4,7)$. Therefore, $\mathbb{P}\mathcal{H}(2,1,1)$ can be compactified by the miniversal deformation space of $(Y,q)$ as the weighted projective space
${\rm Def}(Y,q)\cong\mathbb{P}(1,3,4,6,7,8,9,12).$
Remark 5.3.
The structure of $\mathbb{P}{\rm Def}^{-}(Y,q)$ for monomial singularities with fixed semigroups has also been understood fairly well in low genus; see [Nak08, Ste23].
Moreover, as mentioned in Section 5.14, we can add entries of 0 into the signature $\mu$, i.e., zeros of order 0, to generalize the setting of nonvarying strata. Nevertheless, increasing the number of zero entries can lead to more complicated deformations for the corresponding singularities. For example, if $(Y,q)$ is the elliptic $n$-fold point, then $T^{2}$ is nonzero for $n\geq 6$ and $\mathbb{P}{\rm Def}^{-}(Y,q)$ is not a weighted projective space for $n=6$; see [Ste25, Proposition 1.2] and [LP19, Theorem 1.5.7].
@@ 5.14. Nonvarying strata with ordinary marked points
We can generalize the definition of nonvarying strata by including zeros of order 0, i.e., ordinary marked points. Let $\mu^{\prime}=(m_{1},\ldots,m_{n},0,\ldots,0)$, where $\mu=(m_{1},\ldots,m_{n})$ is one of the above nonvarying signatures and the new parameters $s_{1},\ldots,s_{k}$ correspond to the branches of the last $k$ ordinary points. Up to rescaling the parameters, we can assume that
@@
$y$ $=t_{1}^{3}\oplus 0\oplus 0\cdots\oplus 0,\$
@@
The above observation also leads to an application towards the $K(\pi,1)$-conjecture reviewed in Section 1.3. The following is a refined formulation of Theorem 1.3.
@@
Given $(Y,q)\in S(\mu)$, let $(Y^{\prime},q^{\prime})\in S(\mu,0)$ be the singularity class produced by adding an ordinary marked point away from the zeros $p_{1},\ldots,p_{n}$ in the construction of $(Y,q)$. Then $\mathbb{P}{\rm Def}^{-}_{s}(Y^{\prime},q^{\prime})$ is the universal curve punctured at $p_{1},\ldots,p_{n}$ over $\mathbb{P}{\rm Def}^{-}_{s}(Y,q)$. Moreover, if $\mathbb{P}{\rm Def}^{-}_{s}(Y,q)$ is $K(\pi,1)$, then $\mathbb{P}{\rm Def}^{-}_{s}(Y^{\prime},q^{\prime})$ is also $K(\pi,1)$.
In particular, the $K(\pi,1)$-property holds for the discriminant complement in the miniversal deformation space of the singularities $U_{7}$, $U_{8}$, $U_{9}$, $S_{2g+2}$, and $S_{2g+3}$, which arise from the nonvarying strata for $\mu^{\prime}=(4,0)^{\operatorname{odd}}$, $(3,1,0)$, $(6,0)^{\operatorname{even}}$, $(2g-2,0,0)^{\operatorname{hyp}}$, and $(g-1,g-1,0,0)^{\operatorname{hyp}}$, respectively.
We refer to [Loo84a, (7.22)] and [Ste15, Table 2] for the labeling of the above singularities and their defining ideals.
@@
The first claim follows from the preceding description about how to add the parameter $s$ into the parameterization for the new branch attached to the ordinary marked point.
For the second claim, note that for 2g2+n>02g-2+n>0, a curve of genus $g$ punctured at $n$ points is $K(\pi,1)$. Then the desired claim follows from the standard exact homotopy sequence of fibration if the base is also $K(\pi,1)$.
Finally, note that the $K(\pi,1)$-property holds for $ADE$ singularities by [Bri73] and [Del72], which include those from the nonvarying strata by deleting the entries of 0 from $\mu^{\prime}$. Moreover, the concerned singularities are isolated complete intersections and their Milnor numbers are equal to the dimensions of the corresponding strata of differentials; see [Gre20, Theorem 7.2.22] and Remark 2.4. Therefore, $\mathbb{P}{\rm Def}^{-}_{s}=\mathbb{P}{\rm Def}_{s}$ holds for them, which implies the last claim.
For completeness, we list the defining ideals and parameterizations for these singularities as follows:
$U_{7}\colon$ $(xz-y^{2},yz-x^{3})$
$x=t^{3}\oplus 0,$
$y=t^{4}\oplus 0,$
$z=t^{5}\oplus s.$
$U_{8}\colon$ $(xz-y^{2},yx^{2}-yz)$
$x=t_{1}^{2}\oplus t_{2}\oplus 0,$
$y=t_{1}^{3}\oplus 0\oplus 0,$
$z=t_{1}^{4}\oplus 0\oplus s.$
$U_{9}\colon$ $(xz-y^{2},yz-x^{4})$
$x=t^{3}\oplus 0,$
$y=t^{5}\oplus 0,$
$z=t^{7}\oplus s.$
$S_{2g+2}\colon$ $(xz_{1}-xz_{2},z_{1}z_{2}-x^{2g-1})$
$x=t^{2}\oplus 0\oplus 0,$
$y=t^{2g+1}\oplus 0\oplus 0,$
$z_{1}=t^{2g-1}\oplus s_{1}\oplus 0,$
$z_{2}=t^{2g-1}\oplus 0\oplus s_{2}.$
$S_{2g+3}\colon$ $(xz_{1}-xz_{2},x^{g}z_{2}-z_{1}z_{2})$
$x=t_{1}\oplus t_{2}\oplus 0\oplus 0,$
$y=t_{1}^{g+1}\oplus(-t_{2}^{g+1})\oplus 0\oplus 0,$
$z_{1}=t_{1}^{g}\oplus 0\oplus s_{1}\oplus 0,$
$z_{2}=t_{1}^{g}\oplus 0\oplus 0\oplus s_{2}.$
We remark that for $S_{2g+2}$ and $S_{2g+3}$, the original parameter $y$ in each case can be generated by the other parameters, so $y$ does not appear as a generator in the defining ideal. ∎
5.15. Nonvarying hyperelliptic loci
Besides the hyperelliptic components of the strata, in general, one can define the locus of hyperelliptic differentials $\mathcal{H}(\mu)^{\operatorname{hyp}}$ in $\mathcal{H}(\mu)$ for $\mu=(2a_{1},\ldots,2a_{m},b_{1},b_{1},\ldots,b_{n},b_{n})$, where differentials in $\mathcal{H}(\mu)^{\operatorname{hyp}}$ are defined on hyperelliptic curves such that each zero of order $2a_{i}$ is a Weierstrass point and each pair of zeros of order $b_{j}$ is hyperelliptic conjugate.
Theorem 5.5.
The construction in Section 2.2 produces isomorphic singularities for all hyperelliptic differentials in $\mathcal{H}(\mu)^{\operatorname{hyp}}$, i.e., $S(\mu)^{\operatorname{hyp}}$ consists of a unique singularity class (up to relabeling the marked zeros).
Proof.
Let $t_{i}$ be the parameter of the rational branch associated to the zero of order $a_{i}$, and let $r_{j}$ and $s_{j}$ be the parameters associated to the pair of zeros of order $b_{j}$. By the condition (G4), up to rescaling the parameters, we can assume that all pairwise differences of $t_{i}^{2a_{i}+1}$, $r_{j}^{b_{j}+1}$, and $s_{k}^{b_{k}+1}$ are contained in $R$. Additionally, by the hyperelliptic assumption, we have $t_{i}^{2}\in R$ and $r_{j}+\zeta_{j}s_{j}\in R$ for all $i$ and $j$, where $\zeta_{j}=e^{\pi{\rm i}/(b_{j}+1)}$ such that $(r_{j}+\zeta_{j}s_{j})^{b_{j}+1}=r_{j}^{b_{j}+1}-s_{j}^{b_{j}+1}$.
Note that all even powers of $t_{i}$ and all odd powers of $t_{i}$ with exponents $\geq 2a_{i}+3$ are contained in $R$. Moreover, all powers of $r_{j}$ and $s_{j}$ with exponents $\geq b_{j}+2$ as well as $r_{j}^{c}+(\zeta_{j}s_{j})^{c}$ for all exponents $c$ are contained in $R$. It follows that the sum $|\alpha|$ of the entries in the gap sequence $\alpha_{1},\alpha_{2},\ldots]$ of the singularity is bounded above by
$\sum_{i=1}^{m}a_{i}+\sum_{j=1}^{n}b_{j}+1=g,$
where we used the fact that $\sum_{i=1}^{m}2a_{i}+\sum_{j=1}^{n}2b_{j}=2g-2$ and the last one is from $\alpha_{\max\{2a_{i}+1,b_{j}+1\}}=1$ by (G5). Since $|\alpha|$ has to be $g$, the above functions thus generate $R$, which determines the corresponding singularity uniquely. ∎
Remark 5.6.
Theorem 5.5 implies that $\mathbb{P}\mathcal{H}(\mu)^{\operatorname{hyp}}\subset\mathbb{P}{\rm Def}^{-}_{s}(Y,q)$, where $(Y,q)$ is the unique singularity class in $S(\mu)^{\operatorname{hyp}}$. However, $\mathbb{P}{\rm Def}^{-}_{s}(Y,q)$ may contain non-hyperelliptic differentials, i.e., the preceding containment can be strict in general. For example, $\mathbb{P}\mathcal{H}(2,1,1)^{\operatorname{hyp}}$ is a proper subspace of $\mathbb{P}\mathcal{H}(2,1,1)$, where the latter is a nonvarying stratum and equal to $\mathbb{P}{\rm Def}^{-}_{s}(Y,q)$ by Theorem 1.2.
Additionally, up to a finite choice of ordering the marked zeros, the hyperelliptic locus $\mathbb{P}\mathcal{H}(\mu)^{\operatorname{hyp}}$ can be identified with $\mathcal{M}_{0,2g+2+n}$ by marking the images of the Weierstrass points and the conjugate pairs. Note that $\mathcal{M}_{0,2g+2+n}$ can be interpreted as the configuration space of points in the real plane and hence is $K(\pi,1)$. If we know $\mathbb{P}\mathcal{H}(\mu)^{\operatorname{hyp}}=\mathbb{P}{\rm Def}^{-}_{s}(Y,q)$, then it would imply that the locus of smooth deformations in $\mathbb{P}{\rm Def}^{-}(Y,q)$, viewed as the complement of the discriminant locus of singular deformations, is also $K(\pi,1)$.
It is thus a meaningful question to determine when $\mathbb{P}\mathcal{H}(\mu)^{\operatorname{hyp}}=\mathbb{P}{\rm Def}^{-}_{s}(Y,q)$. We will provide a criterion as follows. Denote by $w_{i}$ the zero of order $2a_{i}$ and $u_{j},v_{j}$ the pair of zeros of order $b_{j}$. Let $C$ be a smooth deformation contained in $\mathbb{P}{\rm Def}^{-}_{s}(Y,q)$. Recall that the divisor at infinity is the $\mathbb{Q}$-Cartier divisor
$C_{\infty}=\frac{1}{\ell}\left(\sum_{i=1}^{m}(2a_{i}+1)w_{i}+\sum_{j=1}^{n}(b_{j}+1)(u_{i}+v_{i})\right),$
where $\ell=\operatorname{lcm}\{2a_{1}+1,\ldots,2a_{m}+1,b_{1}+1,\ldots,b_{n}+1\}$.
Proposition 5.7.
Given $\mu=(2a_{1},\ldots,2a_{m},b_{1},b_{1},\ldots,b_{n},b_{n})$, suppose there exists a value of $k$ such that $\lfloor kC_{\infty}\rfloor$ as a linear combination of $w_{i}$ and $u_{j},v_{j}$ has an even coefficient for each $w_{i}$, has the same coefficients for each pair $u_{j},v_{j}$, and satisfies $0<\deg\lfloor kC_{\infty}\rfloor<2g-2$. Then $\mathbb{P}\mathcal{H}(\mu)^{\operatorname{hyp}}=\mathbb{P}{\rm Def}^{-}_{s}(Y,q)$, where $(Y,q)$ is the unique singularity class in $S(\mu)^{\operatorname{hyp}}$ (modulo the ordering of the marked zeros).
Proof.
By Clifford’s theorem, if there is a divisor $D$ on $C$ such that $0<\deg D<2g-2$ and $h^{0}(C,D)=\frac{1}{2}\deg D+1$, then $C$ is hyperelliptic, and moreover, $D$ is linearly equivalent to an effective linear combination of Weierstrass points and conjugacy pairs, where the coefficients of the Weierstrass points are even.
As discussed in Section 2, note that $h^{0}(C,\lfloor\lambda C_{\infty}\rfloor)$ is the same for all deformations in $\mathbb{P}{\rm Def}^{-}_{s}(Y,q)$, including those hyperelliptic deformations in the subspace $\mathbb{P}\mathcal{H}(\mu)^{\operatorname{hyp}}$. Therefore, taking $D=\lfloor kC_{\infty}\rfloor$ as in the assumption gives a divisor that satisfies the hyperelliptic Clifford’s bound. ∎